OpenStudy (anonymous):

Find the literinization l(x) of f(x) at x=a f(x) = x + 1/x, a=4

OpenStudy (anonymous):

\[f(x)=4+1/4\] \[f(x)=0.25+4\]

OpenStudy (anonymous):

f(4)=0.25

OpenStudy (anonymous):

so you just plug 4 in ? oh. You don't need to find the derivative or anything?

OpenStudy (anonymous):

Just plug in the given number to the equation.